# Inequalities of Brascamp–Lieb type on the Heisenberg group

### 04 Jan 2024 v1

@@ Abstract.
We prove both necessary and sufficient conditions for $L^{p}$-boundedness of certain multilinear generalized Radon transforms that arise as Heisenberg group analogues of the Brascamp–Lieb inequalities on Euclidean space. The necessary and sufficient conditions coincide in some important special cases.
@@ 1. Introduction
$\dim W^{\perp}-\dim V\leq\displaystyle\sum_{j=1}^{m}\frac{\dim W^{\perp}-\dim(L_{j}V^{\perp})^{\perp}}{p_{j}}+\frac{\dim L_{j}(W^{\perp})-\dim V}{p_{j+m}};$ (C1)
$\dim W^{\perp}-\dim V\geq\displaystyle\sum_{j=1}^{m}\frac{\dim W^{\perp}-\dim L_{j}V}{p_{j}}+\frac{\dim(L_{j}W)^{\perp}-\dim V}{p_{j+m}}.$ (C2)
We observe that $\dim(L_{j}V^{\perp})^{\perp}\leq\dim(L_{j}V)$, for all $V$. If equality holds, then, taking $W=V^{\perp}$, we see that (C1) and (C2) are equivalent to
@@
which is stronger than (C1) and (C2) in general.
@@
In Section 2, we will prove that (C) is not necessary in general, though Theorem 1.2 is sharp when the $L_{j}$ are coordinate projections. Determining sharp necessary and sufficient conditions for (1.4) seems like an interesting question and, furthermore, there are other natural generalizations one might consider, such as replacing $L_{j}\oplus\mathbb{I}_{n}$ with a more general linear map, replacing $\mathbb{H}^{n}$ with a more general nilpotent Lie group, or by inserting a cutoff function into the integral defining (1.1), but we do not address these generalizations here.
@@
and Fässler–Pinamonti had established sharp $L^{p}$ bounds in [12]. Though the form (1.1), with maps (1.6), and its sharp or sharp-to-endpoints $L^{p}$ inequalities had implicitly appeared in prior work [15, 18], it had not previously been recognized as a variant of the Loomis–Whitney inequality, and this arrangement naturally raises the question of generalizations in the spirit of the Brascamp–Lieb inequalities. In [5], Bramati used an analogue of the nonlinear heat flow to establish analogous inequalities in a wider class of groups (graded nilpotent Lie algebras), but with a more restrictive condition on the maps involved (the analogue of the case of geometric data from [4]). While some cases that we consider are already in [5], our methods are quite different.
@@ 2. Examples
More precisely, for $1\leq j\leq m$, consider $M_{j}\coloneqq\{m_{j}^{1},\dots,m_{j}^{n_{j}}\}\subset\{1,\dots,n\}$ and $L_{j}$ the orthogonal projections onto subspaces spanned by coordinates whose indices are in $M_{j}$, and set $L_{j+m}:=L_{j}$. Then elementary linear algebraic considerations show that conditions (B1) and (B2) hold for arbitrary $V\leq\mathbb{R}^{n}$ if and only if they hold for all coordinate subspaces. Furthermore, for a coordinate subspace $V$, $L_{j}(V^{\perp})=(L_{j}V)^{\perp}$, $1\leq j\leq m$, so conditions (C1) and (C2) coincide and imply (C) for coordinate subspaces.
In fact, the proof of Theorem 1.2 may be easily modified (with this modification left to the reader) to prove the following.
@@
Given coordinate subspaces $V_{1},\ldots,V_{2m}\leq\mathbb{R}^{n}$, with $V_{j}=V_{j+m}$, inequality (1.4) holds if and only if conditions (A), (B1), (B2), (C) hold for all coordinate subspaces $V\leq\mathbb{R}^{n}$.
@@ 3. Necessity
It is immediate that $|\Omega|\sim R^{\dim V-\dim W^{\perp}}$. Using skew transformations in the $\pi_{j}(\mathbb{H}^{n})$, we also see that for $10<R<$ 1,
$|\pi_{j}(\Omega)|\lesssim\begin{cases}R^{\dim(L_{j}V^{\perp})^{\perp}-\dim W^{\perp}},&1\leq j\leq m\\ R^{\dim V-\dim L_{j}(W^{\perp})},&m+1\leq j\leq 2m.\end{cases}$
Thus, by (3.1),
$1\lesssim R^{C(\mathbf{p},V)},\qquad C(\mathbf{p},V):=\sum_{j=1}^{m}\frac{\dim(L_{j}V^{\perp})^{\perp}-\dim W^{\perp}}{p_{j}}+\frac{\dim V-\dim L_{j}(W^{\perp})}{p_{j+m}}.$
Sending $R\searrow 0$ gives (C1). Similarly, taking $R>1$ and sending $R\nearrow\infty$ yields (C2). ∎
@@ 4. Sufficiency
Now we assume $m>1$. Since we have assumed that equality never holds in (B), we must have $p_{j}\in\{1,\infty\}$ for some $j$ for $\mathbf{p}$ to be extreme. Indeed, if $p_{k}=1$ for some $k$, then by (A) and the fact that $\mathbf{p}\in[1,\infty]^{2m}$, we have $p_{j}=\infty$ for all $j\neq k$. Therefore, $\sum_{j=1}^{2m}1/p_{j}=1/p_{k}=1$, so $\{0\}$ is critical, contradicting the assumption.
Since we have seen that no $p_{j}$ can equal 1, we now assume $p_{j}=\infty$ for $l$ values of $j$. Because (C) may be rewritten as a system of equations for $p_{j}^{-1}-p_{j+m}^{-1}$, (A) and (C) collectively determine at most $m+1$ independent linear equations. Hence, for $\mathbf{p}$ to be extreme, we must have $l\geq 2m-(m+1)=m-1$.
If $l\geq m+1$, then $p_{j}=p_{j+m}=\infty$ for some $1\leq j\leq m$ by pigeonholing, so Conclusion i holds with $j=j^{\prime}$.
If $l=m-1$, (A) and (C) together contain exactly $m+1$ linearly independent equations. In particular, we have exactly $m$ independent equations for $p_{j}^{-1}-p_{j+m}^{-1}$, whence $p_{j}=p_{j+m}$ for all $j$. Since $p_{j}=\infty$ for some $j$, Conclusion i holds with $j=j^{\prime}$.
We turn now to the last case, that $l=m$ and exactly half of the $p_{j}$ are infinite. If $p_{j}=p_{j+m}$ for each $j$, then for some $j$, both exponents must be infinite, and we are in Conclusion i. Therefore (C) contributes exactly $m-1$ independent linear equations.
If $L_{r}=L_{s}$ for some $1\leq r\neq s\leq m$, then the coefficients in (C) of $p_{r}^{-1}-p_{r+m}^{-1}$ and $p_{s}^{-1}-p_{s+m}^{-1}$ are equal for every $V$. Thus, by our hypothesis on (C), there can be at most one such pair. Without loss of generality, we may assume $r=1$. Furthermore, if $L_{1}=L_{s}$, we can form a new $2(m-1)$-linear operator by omitting one of the redundant pairs and a new $(2m-2)$-tuple of exponents $\mathbf{\tilde{p}}$, with entries given by
$\tilde{p}_{1}^{-1}:=p_{1}^{-1}+p_{s}^{-1},\qquad\tilde{p}_{1+m}^{-1}=p_{1+m}^{-1}+p_{s+m}^{-1},\qquad\tilde{p}_{j}^{-1}=p_{j}^{-1},$
for $j\notin\{1,s,1+m,s+m\}$. This new operator and $2(m-1)$-tuple of exponents are in the previous case, that (C) yields exactly $m-1$ independent equations for the $\tilde{p}_{j}^{-1}-\tilde{p}_{j+m}^{-1}$, so we are in Conclusion i for $\mathbf{\tilde{p}}$. If $\tilde{p}_{1}=\tilde{p}_{1+m}=\infty$, we obtain $p_{1}=p_{1+m}=p_{s}=p_{s+m}=\infty$. If $\tilde{p}_{j}=\tilde{p}_{j+m}=\infty$ for some $j\neq 1,s$, then $p_{j}=p_{j+m}=\infty$. Either case leads to Conclusion i for $\mathbf{p}$.
Finally, we may assume that the $L_{j}$ are distinct, so their kernels, $K_{j}:=\ker L_{j}$ are also distinct. Reordering indices if necessary, we may assume that $K_{1}$ is maximal with respect to set containment. Since $K_{1}\cap K_{j}$ is nowhere dense in $K_{1}$ for all $2\leq j\leq m$, by the Baire Category Theorem, $K_{1}\not\subseteq\bigcup_{j=2}^{m}K_{j}$. Pick any $v\in K_{1}\setminus\bigcup_{j=2}^{m}K_{j}$, and let $V:=\langle v\rangle$, the span of $v$. Then $L_{1}|_{V}=0$, while $L_{j}|_{V}$ has rank 1 for $2\leq j\leq m$. Applying (C) with this $V$, $p_{1}=p_{1+m}$. Finally, since $m$ of the $p_{j}$ are infinite, by pigeonholing, this implies the existence of an infinite pair, and we are again in Conclusion i. ∎

### 05 Apr 2024 v2

@@ Abstract.
We prove both necessary and sufficient conditions for $L^{p}$-boundedness of certain multilinear generalized Radon transforms that arise as Heisenberg group analogues of the Brascamp–Lieb inequalities on Euclidean space. The necessary and sufficient conditions coincide in some important special cases, but differ in others.
@@ 1. Introduction
$\dim W^{\perp}-\dim V\geq\displaystyle\sum_{j=1}^{m}\frac{\dim W^{\perp}-\dim L_{j}V}{p_{j}}+\frac{\dim(L_{j}W)^{\perp}-\dim V}{p_{j+m}};$ (C1)
$\dim W^{\perp}-\dim V\geq\displaystyle\sum_{j=1}^{m}\frac{\dim(L_{j}W)^{\perp}-\dim V}{p_{j}}+\frac{\dim W^{\perp}-\dim L_{j}V}{p_{j+m}}.$ (C2)
Here, for any subspace $V\leq\mathbb{R}^{n}$, $V^{\perp}$ is the orthogonal complement of $V$ in $\mathbb{R}^{n}$, while $(L_{j}V)^{\perp}$ is the orthogonal complement of $L_{j}V$ in $L_{j}(\mathbb{R}^{n})=V_{j}$.
We see that $\dim(L_{j}V^{\perp})^{\perp}\leq\dim(L_{j}V)$, and in the case of equality, (C1) and (C2) (with $W=V^{\perp}$) are equivalent to
@@
which is stronger than (C1) and (C2) in the more general case $W\leq V^{\perp}$.
@@
In Section 2, we will prove that (C) is not necessary in general, though Theorem 1.2 is sharp when the $L_{j}$ are coordinate projections whose kernels direct sum to $\mathbb{R}^{n}$. Determining sharp necessary and sufficient conditions for (1.4) seems like an interesting question and, furthermore, there are other natural generalizations one might consider, such as replacing $L_{j}\oplus\mathbb{I}_{n}$ with a more general linear map, replacing $\mathbb{H}^{n}$ with a more general nilpotent Lie group, or by inserting a cutoff function into the integral defining (1.1), but we do not address these generalizations here.
@@
and Fässler–Pinamonti had established sharp $L^{p}$ bounds in [12]. Though the form (1.1), with maps (1.6), and its sharp or sharp-to-endpoints $L^{p}$ inequalities had implicitly appeared in prior work [15, 18], it had not previously been recognized as a variant of the Loomis–Whitney inequality, and this arrangement naturally raises the question of generalizations in the spirit of the Brascamp–Lieb inequalities. Some related questions were posed in [5] in a wider class of groups (graded nilpotent Lie algebras). (See also [12].)
@@ 2. Examples
More precisely, for $1\leq j\leq m$, consider $M_{j}\coloneqq\{m_{j}^{1},\dots,m_{j}^{n_{j}}\}\subset\{1,\dots,n\}$ and $L_{j}$ the orthogonal projections onto subspaces spanned by coordinates whose indices are in $M_{j}$, and set $L_{j+m}:=L_{j}$. Then elementary linear algebraic considerations show that conditions (B1) and (B2) hold for arbitrary $V\leq\mathbb{R}^{n}$ if and only if they hold for all coordinate subspaces. Furthermore, for a coordinate subspace $V$, $L_{j}(V^{\perp})=(L_{j}V)^{\perp}$, $1\leq j\leq m$, so conditions (C1) and (C2) coincide and imply (C) for coordinate subspaces. (However, (C) may fail for subspaces that are not coordinate subspaces.)
By specializing a bit further, we obtain a case in which the necessary and sufficient conditions coincide.
@@
Given coordinate subspaces $V_{1},\ldots,V_{2m}\leq\mathbb{R}^{n}$, with $V_{j}=V_{j+m}$ and $\oplus_{j=1}^{m}\ker L_{j}^{\perp}=\mathbb{R}^{n}$ (recall that $V_{j}=L_{j}(\mathbb{R}^{n})\leq\mathbb{R}^{n}$ and $V_{j}^{\perp}=\ker L_{j}$), inequality (1.4) holds if and only if conditions (A), (B1), (B2), (C) hold for all coordinate subspaces $V\leq\mathbb{R}^{n}$.
(The proof of Theorem 1.2 can easily be modified to prove this result.)
@@ 3. Necessity
It is immediate that $|\Omega|\sim R^{\dim V-\dim W^{\perp}}$. Since $\widehat{L}_{j}x\cdot\widehat{L}_{j}y=x\cdot y-L_{j}x\cdot L_{j}y$, by change of variables, the projections of $\Omega$ have the same volume as the projections via
$\tilde{\pi}_{j}(x,y,t)=\begin{cases}(L_{j}x,y,t+\frac{1}{2}x\cdot y),&1\leq j\leq m,\\ (x,L_{j}y,t-\frac{1}{2}x\cdot y),&m<j\leq 2m.\end{cases}$
Now for $R>1$,
$|\pi_{j}(\Omega)|\lesssim\begin{cases}R^{\dim L_{j}V-\dim W^{\perp}},&1\leq j\leq m\\ R^{\dim V-\dim(L_{j}W)^{\perp}},&m+1\leq j\leq 2m.\end{cases}$
Thus, by (3.1), sending $R\nearrow\infty$ yields (C1); (C2) follows by symmetry. ∎
@@ 4. Sufficiency
Now we assume $m>1$. Because (C) may be rewritten as a system of equations for $p_{j}^{-1}-p_{j+m}^{-1}$, and the second equation in (A) is (C) in the case $V=\mathbb{R}^{n}$, (A) and (C) collectively determine at most $m+1$ independent linear equations. Since we have assumed that $\mathbf{p}$ is extreme and that equality never holds in (B), we must have $p_{j}\in\{1,\infty\}$ for $l\geq 2m-(m+1)=m-1$ values of $j$. However, $p_{j}=1$ cannot hold. Indeed, if $p_{k}=1$ for some $1\leq k\leq 2m$, then by (A) and the fact that $\mathbf{p}\in[1,\infty]^{2m}$, we have $p_{j}=\infty$ for all $1\leq j\neq k\leq 2m$, whence, $\sum_{j=1}^{2m}1/p_{j}=1/p_{k}=1$, so $\{0\}$ is critical, contradicting our assumption that (B) never holds.
Now we have that $p_{j}=\infty$ for $l\geq m-1\geq 1$ values of $j$. If $l\geq m+1$, then $p_{j}=p_{j+m}=\infty$ for some $1\leq j\leq m$ by pigeonholing, so Conclusion i holds with $j=j^{\prime}$. If $l=m-1$, (A) and (C) together contain exactly $m+1$ linearly independent equations. In particular, (C) contains $m$ independent homogeneous equations for $p_{j}^{-1}-p_{j+m}^{-1}$, whence $p_{j}=p_{j+m}$ for all $j$. Since $p_{j}=\infty$ for some $j$, Conclusion i holds with $j=j^{\prime}$.
We turn now to the last case, that $l=m$ and exactly half of the $p_{j}$ are infinite. If $p_{j}=p_{j+m}$ for some $j$, then, by pigeonholing, for some (possibly different) $j^{\prime}$, it must hold that $p_{j^{\prime}}=p_{j^{\prime}+m}=\infty$, so we are in Conclusion i.
Now we may assume that exactly one of $p_{j}$, $p_{j+m}$ equals infinity for every $1\leq j\leq m$. Suppose that the $L_{j}$, $1\leq j\leq m$, are distinct; equivalently, their kernels $K_{j}:=\ker L_{j}=V_{j}^{\perp}$ are distinct. Reordering indices if necessary, we may assume that $K_{1}$ is not contained in any $K_{j}$ with $1<j\leq m$. Therefore $K_{1}\not\subset\bigcup_{j=2}^{m}K_{j}$. Let $v\in K_{1}\setminus\bigcup_{j=2}^{m}K_{j}$, and consider $V:=\langle v\rangle$, the span of $v$. Applying (C) with this $v$, $p_{1}=p_{1+m}$, but this contradicts our assumption that exactly one of each $p_{j},p_{j+m}$ is infinite. Tracing back, the $L_{j}$ cannot all be distinct.
Finally, we may assume that $L_{r}=L_{s}$ for some $1\leq r\neq s\leq m$. The equations (A) and (C) cannot distinguish $p_{r}^{-1}$ from $p_{s}^{-1}$, so for $\mathbf{p}$ to be extreme in the absence of (B), it must hold that $p_{r}=\infty$ or $p_{s}=\infty$ (since neither can equal 1). Similarly, it must also hold that $p_{r+m}=\infty$ or $p_{s+m}=\infty$. This once again puts us in the situation described by Conclusion i, completing the proof of the Lemma. ∎